OpenStudy (jazzyluvsu123):

X * 1/4=1/3 How do I get my answer??

OpenStudy (jazzyluvsu123):

@mathmale

pooja195 (pooja195):

Use cross multiplcation \[\huge~\rm~\frac{ (x)(1) }{ 4 }=\frac{ 1 }{ 3 }\]

OpenStudy (jazzyluvsu123):

I thought you Divided it ?

OpenStudy (jazzyluvsu123):

So my answer would be 1/12??

pooja195 (pooja195):

How did you get that?

OpenStudy (jazzyluvsu123):

Cross multiply 1*1 is 1 and 3*4 is 12

OpenStudy (jazzyluvsu123):

so 1/12

pooja195 (pooja195):

That's incorrect. @Ace8407 TEACH the user do not guide them to an online calculator.

pooja195 (pooja195):

3 * 1x=3x 3x=4 now isolate x divide both sides by 3 what do you get? x=?

OpenStudy (jazzyluvsu123):

Ooo @pooja195 i went directly across instead of Crossing it i understand it now xD sorry

pooja195 (pooja195):

yw ^_^

OpenStudy (jazzyluvsu123):

4*1 is 4 and 3*1 = 4/3

pooja195 (pooja195):

Well done :-)

OpenStudy (jazzyluvsu123):

Thanks ;]

pooja195 (pooja195):

yw ^_^